OpenStudy (anonymous):

find the limit as x -->0 (sin(kx))/(kx^2-2kx)

OpenStudy (jamesj):

Do you know the limit as x --> 0 of sin(x) / x ?

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

noooo its 1!

OpenStudy (jamesj):

No, it's 1. Using this result, you can arrange your function in such a way as you can see the result.

OpenStudy (anonymous):

how do i rearrange the function?

OpenStudy (jamesj):

\[\frac{ \sin(kx) }{kx^2-2kx} = \frac{ \sin(kx)}{kx} . \frac{1}{x - 2}\]

OpenStudy (anonymous):

take the derivative of the nominator and denominator: kcos(kx)/(2kx -2k) plug x = 0 k / -2k = -1/2

OpenStudy (anonymous):

i thought the answer was 1?

OpenStudy (jamesj):

Now \[\lim_{x \rightarrow 0} \frac{\sin kx}{kx} = 1\] Hence your limit is ...

OpenStudy (anonymous):

that one guy said the limit was (-1/2) though?

OpenStudy (jamesj):

He was talking about the entire problem.

OpenStudy (jamesj):

and he's right, but he using a rule for limits that you may not know yet.

OpenStudy (anonymous):

no no i remember that rule now but thank you

OpenStudy (jamesj):

For the record my argument is elementary: \[\lim_{x \rightarrow 0} \frac{\sin kx}{kx^2 - 2kx} = \lim_{x \rightarrow 0} \frac{\sin kx}{kx}.\lim_{x \rightarrow 0} \frac{1}{x-2} \] which is equal to 1 times -1/2 I.e., -1/2.

OpenStudy (anonymous):

ok cool deal